Crew 218 Sol Summary 29Dec2019

[title Sol Summary – December 29th]

[category sol-summary]

Sol: 07

Summary Title: Breath of Fresh Martian Air

Author’s name: Pat Pesa

Mission Status: Crew is excited about Today’s progress and the clear weather ahead.

Sol Activity Summary: Crew 218 Finally got out of the Habitat for an EVA this afternoon! Along with continued work on our Indoor/Outdoor Climate monitoring, we started our Journalist’s project in Detecting Radio Signal Strength in the local area.

Look ahead plan: Hoping for more EVAs for tomorrow after more melting occurs. Tomorrow we will also try to finish crew interviews and training videos.

Anomalies in work: None

Weather: Partly Cloudy, around 25 degrees

Crew physical status: Healthy

EVA: none

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, EVA request, science report

Support Requested: none

Crew 218 Science Report 29Dec2019

Crew 218 Science Report 29-DEC-19
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

No EVA: nothing to report.

2. Mars surface weather:

No EVA: nothing to report.

3. Subsurface structure on Mars:

No EVA: nothing to report.

4. Detecting radio signal strength:

No EVA: nothing to report.

5. EVA workload analysis:

No EVA: nothing to report.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

No EVA: nothing to report.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

Nothing to report.

8. Messier and other space objects for outreach:

Cloudy Weather: nothing to report.

9. Reliability and maintenance:

No EVA: nothing to report.

10. Medical readings in preparation for future crew-wide project:

The Commander, Dr. Cesare Guariniello, has donned the Zephyr bio-medical sensor at 11:00AM, and is testing real-time measurement infrastructure for EVA and habitat medical monitoring. He will wear the instrument continuously for 24-48 hours.

11. Collection of clay, shale, and hematite samples:

No EVA: nothing to report.

12. Media and outreach:

Nothing to report.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

Crew 218 GreenHab Report 29Dec2019

Crew 218 GreenHab Report 29-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 23°C; 18%

11:00AM

Floor Unit: 18°C

Electronic: 24°C

humidity 18%

12:35PM

Floor Unit: 18°C

Electronic: 23.7°C

humidity 18%

3:15PM

Floor Unit: 17°C

Electronic: 19.8°C

humidity 20%

5:45PM

Floor Unit: 14°C

Electronic: 19.2°C

humidity 20%

Max: 28.8; 21%

Min: 17.2°C; 16%

Hours of supplemental light: Light system 6:15PM-11:15AM.

Daily water usage for crops: 5.4Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~100 Gal.

Time(s) of watering for crops:

6:00PM

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: Zucchini, Cherry Tomato, and Pepper had first sprouts.

—Door was opened while in GreenHab, but closed upon leaving.

Harvest:

Aloe Vera 9g

EVA Report 29Dec2019

EVA #: 3

Author: Cesare Guariniello

Purpose of EVA: weather observation project; radio mapping project; collection of geological samples

Start time: 12:00

End time: 13:18 (including engineering EVA)

Narrative: one walking EVA has been approved for today. Due to the presence of snow, the commander joined the EVA for better knowledge of the terrain. Working on snow was easy and safe, and a couple of narrow streams formed by the melted water were less than one step across. At the end of the simulated depressurization cycle, the crew noticed that the front door was still blocked by some snow and Pat completed the shoveling process. At the same time, both Ben and Luz Ma experienced problems with their equipment, therefore a cycle of re-pressurization was initiated and the crew re-entered the habitat to fix a hat and a pair of shoes. After the second depressurization, Cesare, Luz Ma and Ben egressed and Luz Ma and Ben became their measurements with the help of Cesare. Then, the crew grabbed their bag with tools, EVA first aid kit, map, and rock hammer and walked towards the MDRS sign and out along Cow Dung Road to Marble Ritual. Radio measurements were collected along the way. Once at destination, both weather data and radio measurements were collected before the crew headed back to the habitat. An engineering EVA was performed before ingress.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): E518700, N4250800

EVA Participants: LuzMa Agudelo, ENG; Ben Durkee, JOU; Cesare Guariniello, CMD.

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Walking

Operations Report – December 29th

Crew 218 Operations Report 29-Dec-19

SOL: 07
Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA
Generator: run
Hours run: 11
From what time last night: 2000
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before the generator is run at night: 78%
Diesel Reading – 77%
Station Propane Reading – 71%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 21
Water Meter (units): 0147111.1
Water (static tank) (gal): 395
Static to Loft Pump used – No
Water in Green Hab (gal): 100
Water in Science Dome (gal): 0
Toilet tank emptied: No

Deimos rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Not used
Hours: 119.2
Beginning charge: (Before EVA):
Ending charge: (On return from EVA, before recharging):
Currently charging: Yes
Opportunity rover used: Still in the workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: Not used
Hours: 124.4
Beginning charge:
Ending charge: (On return from EVA, before charging):
Currently charging: Yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A
HabCar used and why, where? No
CrewCar used and why, where? No
General notes and comments: N/A
Summary of the internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: The pipe from the static tank through the loft tank froze before we did the regular daily filling yesterday (12/28/2019). We tried to fill the tank several times during the day today without luck. The pipes are still frozen, we are rationing the water, and we will keep checking on the status of the pipe.

Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal
Questions, concerns, and requests to Mission Support: NA

Journalist Report – December 28th

Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 06

I swear these morning workouts get more and more intense every day. Sitting up has now become an ordeal, and don’t even get me started on descending our glorified ladder of a staircase. I feel like a 21-year-old going on octogenarian – perhaps we age faster on Mars? I may have to request some dentures in the next cargo resupply craft sent from Earth…

In spite of our body aches, we spent a good chunk of time today sweeping the snow from our inter-Habitat tunnels again. We still haven’t been able to locate that pesky leak that’s allowing the precipitation in. I suspect it is the Martian mouse from Sol 04 trying to take us out one by one. We haven’t been able to track him down, yet, but have affectionately named him Marvin.

It’s very interesting the way the mind wanders when given the time and space to do so.

“When Dwayne ‘The Rock’ Johnson studies his family history, is it considered genealogy or geology?”

This is the question Pat (the crew geologist) and I pondered over lunch. After involving the rest of the crew in our whirlwind debate, the general consensus became the latter, but we welcome outside opinions on the matter.

After discussing our philosophical quandary, the curtain of clouds parted for the first time in days to reveal unfettered, unfiltered sunlight. Not only does this imply the most snow melting power we’ve had yet, but it also means the sun was finally ripe for observation. Cesare rushed out to the observatory to carpe this rare diem.

Once he had methodically set up all of the instrumentation, he invited us in individually to join in his observation. Eventually my turn rolled around, and I removed my glasses and peered into the gauntlet of mirrors and lenses. We fiddled with the knobs until I got a view of the sun. Circular, as expected. A bit more adjustment yielded better focus, and I was able to watch the surface of the sun subtly pulsate and squirm, like a colossal heart of nuclear fusion.

Cesare spent some time adjusting machinery and software far beyond in an attempt to capture some solar activity. We weren’t able to observe any sun spots or granulation, but in hindsight that may be a blessing considering Mars’s lack of shielding against such radioactive tumult.

The sun soon escaped our grasp and retreated behind the horizon, so we too retreated to the Hab. A dash of Smash Bros. to wind down, a pinch of pepper on our dinner, and the evening had escaped us in a similar fashion. We’ve a cold night ahead of us, but hopefully the sun will return unobstructed to thaw our Martian landscape. If the stars align, we may even be able to go on EVA tomorrow and launch our scientific discoveries into full swing. Fingers crossed!

Science Report – December 28th

Crew 218 Science Report 28-DEC-2019
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

No EVA: nothing to report.

2. Mars surface weather:

No EVA: nothing to report.

3. Subsurface structure on Mars:

No EVA: nothing to report.

4. Detecting radio signal strength:

No EVA: nothing to report.

5. EVA workload analysis:

No EVA: nothing to report.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

No EVA: nothing to report.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

The Crew Engineer, Luz Maria Agudelo Urrego, reported that overnight data collection was completed.

We present preliminary results. The x-axis is time in a 24 hour period, the y-axis is temperature (°C), and each color represents a different day in the LuzMa’s room. The threshold of 26.6°C is considered the lower limit for human thermal heat stress.
Unexpectedly, the nighttime temperatures are surpassing the threshold, and are sustained throughout the night.

8. Messier and other space objects for outreach:

Cloudy Weather: nothing to report.

9. Reliability and maintenance:

No EVA: nothing to report.

10. Medical readings in preparation for future crew-wide project:

No EVA: nothing to report.

11. Collection of clay, shale, and hematite samples:

No EVA: nothing to report.

12. Media and outreach:

The Crew Journalist, Benjamin Durkee, has developed interview questions and will interview each crew member.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

DataCollection.pdf

Operations Report – December 28th

Crew 218 Operations Report 28-Dec-19

SOL: 06
Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA
Generator: run
Hours run: 13
From what time last night: 1800
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A Solar— SOC % (Before generator is run at night: 78%
Diesel Reading – 81%
Station Propane Reading – 74%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 35
Water Meter (units): 0147092.6
Water (static tank) (gal): 395
Static to Loft Pump used – Yes
Water in Green Hab (gal): 105
Water in Science Dome (gal): 0
Toilet tank emptied: No

Deimos rover used: Still in workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Not used
Hours: 119.2
Beginning charge: (Before EVA):
Ending charge: (On return from EVA, before recharging):
Currently charging: Yes
Opportunity rover used: Still in workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: Not used
Hours: 124.4
Beginning charge:
Ending charge: (On return from EVA, before charging):
Currently charging: Yes
Notes on rovers: Opportunity and Deimos off-site for maintenance. ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A
HabCar used and why, where? No
CrewCar used and why, where? No
General notes and comments: N/A
Summary of internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: Nominal

Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal
Questions, concerns and requests to Mission Support: NA

LuzMa

ENG

Crew 218

GreenHab Report – December 28th

Crew 218 GreenHab Report 28-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 23.2°C; 19%

11:15AM

Floor Unit: 17°C

Electronic: 21.6°C

humidity 19%

2:50PM

Floor Unit: 19°C

Electronic: 26.0°C

humidity 17%

6:45PM

Floor Unit: 16°C

Electronic: 22.0°C

humidity 19%

Max: 28.5°C; 23%

Min: 17.9°C; 16%

Hours of supplemental light: Light system 7:20PM-12:20AM.

Daily water usage for crops: 5.4Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~105 Gal.

Time(s) of watering for crops:

6:45PM

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: 2 new snow pea sprouts.

—2 snow peas completely dried out post relocation (they were drying out before). Hopeful that they will rehydrate.

—Door closed.

Harvest:

Marjoram: Total: 21.82g

Leaves: 16.08g

Stems: 5.74g

Cilantro: 4.55g